OpenStudy (anonymous):

Determine the number of solutions based on the graph for the following system of equations 2x+5y=7 10y=-4x+14

OpenStudy (anonymous):

go to your graphing calculator. type in y= -2/5X + 7/5 and then type in y= -4/10X + 14/10 and graph it.

OpenStudy (anonymous):

i dont have a graphing calculator o.o

OpenStudy (anonymous):

thats weird. hold on. i dont have my calculator but maybe i can figure it out

OpenStudy (anonymous):

ok lol thanks

OpenStudy (anonymous):

those equations both equal each other so they have all the same points on the graph

OpenStudy (anonymous):

They are infinitely many solutions, explain why?